# Optimal Online Change Detection via Random Fourier Features

### v1

@@ A.3.2 Distribution-free bound
In this section, we show the change detection performance of the proposed Online RFF-MMD algorithm if no pre-change sample is used to estimate the threshold. Instead, we use Theorem 1 to compute the distribution-free threshold sequence $\{\lambda_{n}\mid n\in\mathbb{N}\}$ for a given target ARL. To obtain an EDD estimate, we sample and process $512$ observations from MNIST digit 00 (pre-change) and 1.0241.0241{.}0241.024 samples from digits 1–9 (post-change), respectively, averaging the detection delay over $100$ repetitions. The results are in Figure 6. When comparing to Figure 3 and Figure 5, the figure shows that our method has an increased detection delay, which is due to the looser distribution-free bound (see Section A.3.3 for a numerical comparison). Still, except for the change to the digit $5$ with a guaranteed ARL of $10^{5}$, Online RFF-MMD detects all changes reliably.
@@
A.3.3 Threshold comparison
In this section, we compare the tightness of our thresholds in the offline two-sample testing setting. Specifically, we fix the level $\alpha=0.01$ and let $\mathbb{P}=\mathbb{Q}=\mathcal{N}(0,1)$. We then approximate the $1-\alpha$ quantile of $\operatorname{MMD}_{\hat{K}}\!\left(\hat{\mathbb{P}}_{n},\hat{\mathbb{Q}}_{n}\right)$ (with $n=1{.}000$, $\hat{K}$ approximating the Gaussian kernel with $r=1{.}000$ RFFs, and $\gamma>0$ set by the median heuristic) by (i) obtaining new samples from $\mathbb{P},\mathbb{Q}$ and (ii) permuting a fixed sample from $\mathbb{P},\mathbb{Q}$ for 1.0001.0001{.}0001.000 rounds. Figure 7 shows the respective histograms and the estimated quantiles along with the thresholds obtained by Lemma 5 and Lemma 5, respectively. As one expects, the figure shows that the variance estimate used in Lemma 5 allows to obtain a tighter bound, where we consider the resampling/permutation-based thresholds as ground truth. We emphasize that independent of the threshold used, the resulting test is consistent against any fixed alternative.
@@ A.4 Proofs
This section is dedicated to the proofs of our results stated in the main text. The proof of Theorem 1 is in Appendix A.4.1, that of Theorem 2 is in Appendix A.4.2, that of Theorem 3 is in Appendix A.4.3, and that of our minimax result (Theorem 4) is in Section A.4.4.
@@ A.4.2 Proof of Theorem 2
Write $\pi_{n}=\sqrt{2\left(\log(n/\alpha)+2\log\left(\log_{2}(n)\right)+\log\left(\log_{2}(2n)\right)\right)}$ for each $n\in\mathbb{N}$. Applying standard peeling arguments [59, 53], we have that
@@ A.4.3 Proof of Theorem 3
For the penultimate term in (20b), applying the bound
$\mathbb{E}\left|\operatorname{MMD}_{K}\left[X_{1:n},Y_{1:m}\right]-\operatorname{MMD}_{K}\left[\mathbb{P},\mathbb{Q}\right]\right|\leq 2\left(\frac{1}{\sqrt{n}}+\frac{1}{\sqrt{m}}\right)$
for $X$’s sampled from $\mathbb{P}$ and $Y$’s sampled from $\mathbb{Q}$, whose proof can be found for instance in Section A.2 of [13], together with the bound $\sqrt{x+y}\geq\left(\sqrt{2}/2\right)\left(\sqrt{x}+\sqrt{y}\right)$ for all $x,y\geq 0$, which holds due to the convexity of the square root, one has that
@@
Note that we assume $0\leq K(\cdot,\cdot)\leq 1$. Therefore, arguing as in Lemma 4, one can show that MMD as defined in (6) has the self bounding property with constants (35). Hence, applying Theorem 5 to terms (24a) and (24b) one arrives at (19). Turning to the problem of interest we first make explicit the constants in Theorem 3:
@@
$C_{2}=\frac{1}{3}\left[\left(\frac{\sqrt{2}-1}{4}-\frac{\sqrt{50}+\sqrt{6}}{\sqrt{C_{3}}}\right)\right]$
$C_{3}=\left(\sqrt{6}+\sqrt{50}+\sqrt{54}\right)^{2}/\left(4/(\sqrt{2}-1)\right)^{2}.$
@@
$t_{k^{*}}=\min\left\{t=2^{j}\mid j\in\mathbb{N}\text{ and }t\leq k^{*}\right\}.$ (25b)
Note that condition (11) guarantees that (25a) exists and it can be checked that $C_{2}>0$. Consequently, using (19) and the fact that $k^{*}/2\leq t_{k^{*}}\leq k^{*}$, we obtain that
@@
$\leq\mathbb{P}\Bigg{(}\bigvee_{j=0}^{\left\lfloor\log_{2}\left(\eta+k^{*}\right)\right\rfloor-1}\sqrt{\frac{2^{j}(\eta+k^{*}-2^{j})}{\eta+k^{*}}}\operatorname{MMD}_{\hat{K}}\!\left[X_{A_{k^{*}}},X_{B_{k^{*}}}\right]\leq\sqrt{2}+\lambda_{\eta+k^{*}}\Bigg{)}$
$\leq\mathbb{P}\left(\sqrt{\frac{t_{k^{*}}(\eta+k^{*}-t_{k^{*}})}{\eta+k^{*}}}\operatorname{MMD}_{\hat{K}}\left[X_{1:(\eta+k^{*}-t_{k^{*}})},X_{(\eta+k^{*}-t_{k^{*}}+1):(\eta+k^{*})}\right]>\sqrt{2}+\lambda_{2\eta}\right)$
@@
$\subseteq\left\{\sqrt{k}\times\operatorname{MMD}_{K}\left[\mathbb{P},\mathbb{Q}\right]\geq\sqrt{C_{2}\log\left(\frac{2\eta}{\alpha}\right)}\right\}$
@@ A.4.4 Proof of Theorem 4
Let $\bm{1}=\left(1,\dots,1\right)^{\mkern-1.5mu\mathsf{T}}\in\mathbb{R}^{d}$ and $\bm{0}=\left(0,\dots,0\right)^{\mkern-1.5mu\mathsf{T}}\in\mathbb{R}^{d}$, let $\delta_{\bm{x}}$ denote the Dirac measure (for any set $A\in\sigma\!\left(\mathbb{R}^{d}\right)$ and any $\bm{x}\in\mathbb{R}^{d}$, $\delta_{\bm{x}}\!\left(A\right)=1$ if $\bm{x}\in A$ and 00 otherwise), and let
=$\mathcal{M}^{*}=\left\{\mathbb{P},\mathbb{Q}\in\mathcal{M}_{1}^{+}\left(\mathbb{R}^{d}\right)\mid\mathbb{P}=p\delta_{\bm{1}}+\left(1-p\right)\delta_{\bm{0}},\mathbb{Q}=q\delta_{\bm{1}}+\left(1-q\right)\delta_{\bm{0}},\right.$
$\left.p,q\in\left[1/4,3/4\right]\text{ and }q-p\geq 1/4\right\}.$
@@
$\leq\left(q-p\right)\left[\frac{q}{p}-\frac{1-p}{2-(p+q)}\right]$ (29a)
@@
Therefore, for a putting $C_{K}=2(3/17)\left(K\left(\bm{0},\bm{0}\right)-K\left(\bm{1},\bm{0}\right)\right)$ for the constant in (13), using (30) along with the fact that $\mathcal{M}^{*}\subset\mathcal{M}_{1}^{+}\left(\mathbb{R}^{d}\right)$, we have that
@@
Consequently the theorem is proved if we can find absolute constant $\alpha_{0},\beta_{0}\in(0,1)$ and pre- and post-change distributions $\mathbb{P},\mathbb{Q}\in\mathcal{M^{*}}$ such that for all $\alpha\leq\alpha_{0}$ it holds that
@@
To show (32), one can use a change of measure argument originally due to Lai [27]. In fact one can directly use the version of Lai’s argument adapted to finite sample analysis by Yu et al. [59, Proposition 4.1]. For clarity of exposition, we repeat the argument below. The following holds for arbitrary $\mathbb{P},\mathbb{Q}\in\mathcal{M^{*}}$. For each $n\in\mathbb{N}$ let $\mathcal{F}_{n}$ be the $\sigma$-field generated by $\{X_{i}\}_{i=1}^{n}$ and let $\mathbb{P}^{\otimes n}$ be the restriction of the joint law to $\mathcal{F}_{n}$. We can write
@@ A.4.5 Proof of Lemma 1
Without loss of generality assume that $m\geq n$. Let $\hat{z}_{K}(\cdot)$ be as defined in (8). We have that
$\operatorname{MMD}_{\hat{K}}\left[X_{1:n},Y_{1:m}\right]=\left\|\frac{1}{n}\sum_{i=1}^{n}\hat{z}_{K}\left(X_{i}\right)-\frac{1}{m}\sum_{j=1}^{m}\frac{1}{r}\hat{z}_{K}\left(Y_{j}\right)\right\|_{2}.$
Let $I_{1},\dots,I_{n}$ be mutually disjoint sets satisfying $\cup_{i=1}^{n}I_{i}=\left\{1,\dots,m\right\}$ and $\left\lfloor m/n\right\rfloor\leq\left|I_{i}\right|\leq\left\lfloor m/n\right\rfloor+1$ for each $i=1,...,n$. With $\hat{\mu}_{K}:=\mathbb{E}_{X\sim\mathbb{P}}[\hat{z}_{K}\left(X\right)\mid\omega_{1},\dots,\omega_{r}]$ introduce the quantities
$\tilde{\eta}^{X}_{i}=\hat{z}_{K}\left(X_{i}\right)-\hat{\mu}_{K}\hskip 10.00002pt\text{ for }i=1,\dots,n$
$\tilde{\eta}^{Y}_{j}=\hat{z}_{K}\left(Y_{j}\right)-\hat{\mu}_{K}\hskip 10.00002pt\text{ for }j=1,\dots,m$
$\tilde{\eta}^{X,Y}_{i}=\tilde{\eta}_{i}^{X}-\left\lfloor\frac{m}{n}\right\rfloor^{-1}\sum_{j\in I_{i}}\tilde{\eta}^{Y}_{j}\hskip 10.00002pt\text{ for }i=1,\dots,n.$
It is easy to see that $\operatorname{MMD}_{\hat{K}}\left[X_{1:n},Y_{1:m}\right]=\left\|\frac{1}{n}\sum_{i=1}^{n}\tilde{\eta}^{X,Y}_{i}\right\|_{2}$, and that conditional on the $\omega$’s the $\tilde{\eta}^{X,Y}$’s are independently distributed zero mean random vectors which satisfy the Bernstein condition
$\mathbb{E}\left[\left\|\tilde{\eta}^{X,Y}_{i}\right\|_{2}^{p}\mid\omega_{1},\dots,\omega_{r}\right]$ $\leq\left(2\sqrt{2}\right)^{p-2}\mathbb{E}\left[\left\|\tilde{\eta}^{X,Y}_{i}\right\|_{2}^{2}\mid\omega_{1},\dots,\omega_{r}\right]$
$\leq\frac{1}{2}p!\left(2\sqrt{2}\right)^{p-2}\tilde{\sigma}^{2}\left(\omega_{1},\dots,\omega_{r}\right),\hskip 20.00003pt\text{ for }p\geq 2\text{ and }i=1,\dots,n,$
where, with $i^{\prime}\in\operatorname*{arg\,max}_{i=1,\dots,n}\left|I_{i}\right|$, we let
$\tilde{\sigma}^{2}\left(\omega_{1},\dots,\omega_{r}\right)=\mathbb{E}\left[\left\|\tilde{\eta}^{X,Y}_{i^{\prime}}\right\|_{2}^{2}\mid\omega_{1},\dots,\omega_{r}\right].$
Therefore, applying Theorem 7, we obtain that
$\mathbb{P}\left(\operatorname{MMD}_{\hat{K}}\left[X_{1:n},Y_{1:m}\right]>\varepsilon\right)$
$=\int\dots\int\mathbb{P}\left(\operatorname{MMD}_{\hat{K}}\left[X_{1:n},Y_{1:m}\right]>\varepsilon\mid\omega_{1},\dots,\omega_{r}\right)\mathrm{d}\Lambda\left(\omega_{1}\right)\dots\mathrm{d}\Lambda\left(\omega_{r}\right)$
$\leq\int\dots\int 2\exp\left(-\frac{1}{2}n\varepsilon^{2}\left[\tilde{\sigma}^{2}\left(\omega_{1},\dots,\omega_{r}\right)+2\sqrt{2}\varepsilon\right]^{-1}\right)\mathrm{d}\Lambda\left(\omega_{1}\right)\dots\mathrm{d}\Lambda\left(\omega_{r}\right)$
@@
$=2\mathbb{E}\left[\exp\left(-\frac{1}{2}n\varepsilon^{2}\left[\tilde{\sigma}^{2}\left(\omega_{1},\dots,\omega_{r}\right)+2\sqrt{2}\varepsilon\right]^{-1}\right)\right]$

### v2

@@ A.3.4 Distribution-free bound
In this section, we show the change detection performance of the proposed Online RFF-MMD algorithm if no pre-change sample is used to estimate the threshold. Instead, we use Theorem 1 to compute the distribution-free threshold sequence $\{\lambda_{n}\mid n\in\mathbb{N}\}$ for a given target ARL. To obtain an EDD estimate, we sample and process $512$ observations from MNIST digit 0 (pre-change) and 1 0241\,024 samples from digits 1–9 (post-change), respectively, averaging the detection delay over $100$ repetitions. The results are in Figure 7. When comparing to Figure 4 and Figure 5, the figure shows that our method has an increased detection delay, which is due to the looser distribution-free bound (see Section A.3.5 for a numerical comparison). Still, except for the change to the digit $5$ with a guaranteed ARL of $10^{5}$, Online RFF-MMD detects all changes reliably.
@@
A.3.5 Threshold comparison
In this section, we compare the tightness of our thresholds in the offline two-sample testing setting. Specifically, we fix the level $\alpha=0.01$ and let $\mathbb{P}=\mathbb{Q}=\mathcal{N}(0,1)$. We then approximate the $1-\alpha$ quantile of $\operatorname{MMD}_{\hat{K}}\!\left(\hat{\mathbb{P}}_{n},\hat{\mathbb{Q}}_{n}\right)$ (with n=1 000n=1\,000, $\hat{K}$ approximating the Gaussian kernel with r=1 000r=1\,000 RFFs, and $\gamma>0$ set by the median heuristic) by (i) obtaining new samples from $\mathbb{P},\mathbb{Q}$ and (ii) permuting a fixed sample from $\mathbb{P},\mathbb{Q}$ for 1 0001\,000 rounds. Figure 8 shows the respective histograms and the estimated quantiles along with the thresholds obtained by Lemma 1 and Lemma 5, respectively. As one expects, the figure shows that the variance estimate used in Lemma 1 allows to obtain a tighter bound, where we consider the resampling/permutation-based thresholds as ground truth. We emphasize that independent of the threshold used, the resulting test is consistent against any fixed alternative.
@@ A.4 Proofs
This section is dedicated to our proofs. The proof of Theorem 1 is in Appendix A.4.1, that of Theorem 2 is in Appendix A.4.2, that of Theorem 3 is in Appendix A.4.3, and that of our minimax result (Theorem 4) is in Section A.4.4. The tighter threshold detailed in Lemma 1 is proved in Appendix A.4.5 and we state the proof for our multiple change point detection result (Theorem 5) in Appendix A.4.6.
@@ A.4.2 Proof of Theorem 2
Write
$\pi_{n}=\sqrt{2\left(\log(n/\alpha)+2\log\left(\log_{2}(n)\right)+\log\left(\log_{2}(2n)\right)\right)}$ (17)
for each $n\in\mathbb{N}$. Applying standard peeling arguments [64, 57], we have that
@@ A.4.3 Proof of Theorem 3
For the penultimate term in (20c), applying the bound
$\mathbb{E}\left|\operatorname{MMD}_{K}\left[X_{1:n},Y_{1:m}\right]-\operatorname{MMD}_{K}\left[\mathbb{P},\mathbb{Q}\right]\right|\leq 2\left(\frac{1}{\sqrt{n}}+\frac{1}{\sqrt{m}}\right)$ (22)
for $X$’s sampled from $\mathbb{P}$ and $Y$’s sampled from $\mathbb{Q}$, whose proof can be found for instance in Section A.2 of [14], together with the bound $\sqrt{x+y}\geq\left(\sqrt{2}/2\right)\left(\sqrt{x}+\sqrt{y}\right)$ for all $x,y\geq 0$, which holds due to the concavity of the square root, one has that
@@
Note that we assume $0\leq K(\cdot,\cdot)\leq 1$. Therefore, arguing as in Lemma 4, one can show that MMD as defined in (5) has the bounded differences property with constants (41). Hence, applying Theorem 6 to terms (25a) and (25b) one arrives at (19). Turning to the problem of interest we first make explicit the constants in Theorem 3:
@@
$C_{2}^{-1}=\frac{1}{3}\left[\left(\frac{\sqrt{2}-1}{4}-\frac{\sqrt{50}+\sqrt{6}}{\sqrt{C_{3}}}\right)\right]$
$C_{3}=\left(\sqrt{6}+\sqrt{50}+\sqrt{54}\right)^{2}\times\left(4/(\sqrt{2}-1)\right)^{2}.$
@@
[…]
Note that condition (9) guarantees that (26a) exists and it can be checked that $C_{2}>0$. Let $j^{*}$ be the $j$ appearing in (26b). Consequently, using (19) and the fact that $k^{*}/2\leq t_{k^{*}}\leq k^{*}$, we obtain that
@@
$\leq\mathbb{P}\Bigg(\bigcap_{j=0}^{\left\lfloor\log_{2}\left(\eta+k^{*}\right)\right\rfloor-1}\sqrt{\frac{2^{j}(\eta+k^{*}-2^{j})}{\eta+k^{*}}}\operatorname{MMD}_{\hat{K}}\!\left[X_{I_{j^{*},k^{*}}},X_{J_{j^{*},k^{*}}}\right]\leq\sqrt{2}+\lambda_{\eta+k^{*}}\Bigg)$
$\leq\mathbb{P}\left(\sqrt{\frac{t_{k^{*}}(\eta+k^{*}-t_{k^{*}})}{\eta+k^{*}}}\operatorname{MMD}_{\hat{K}}\left[X_{I_{j^{*},k^{*}}},X_{J_{j^{*},k^{*}}}\right]\leq\sqrt{2}+\lambda_{2\eta}\right)$
@@
$\subseteq\left\{\sqrt{k}\times\operatorname{MMD}_{K}\left[\mathbb{P},\mathbb{Q}\right]\geq\sqrt{C_{3}\log\left(\frac{2\eta}{\alpha}\right)}\right\}$
@@ A.4.4 Proof of Theorem 4
Let $\bm{1}=\left(1,\dots,1\right)^{\mkern-1.5mu\mathsf{T}}\in\mathbb{R}^{d}$ and $\bm{0}=\left(0,\dots,0\right)^{\mkern-1.5mu\mathsf{T}}\in\mathbb{R}^{d}$, let $\delta_{\bm{x}}$ denote the Dirac measure (for any set $A\in\sigma\!\left(\mathbb{R}^{d}\right)$ and any $\bm{x}\in\mathbb{R}^{d}$, $\delta_{\bm{x}}\!\left(A\right)=1$ if $\bm{x}\in A$ and 0 otherwise), and let
$\mathcal{M}^{*}=\left\{\mathbb{P},\mathbb{Q}\in\mathcal{M}_{1}^{+}\left(\mathbb{R}^{d}\right)\mid\mathbb{P}=p\delta_{\bm{1}}+\left(1-p\right)\delta_{\bm{0}},\mathbb{Q}=q\delta_{\bm{1}}+\left(1-q\right)\delta_{\bm{0}},\right$.
[…]
@@
$\leq\left(q-p\right)\left[\frac{q}{p}-\frac{1-q}{2-(p+q)}\right]$ (30a)
@@
Therefore, putting $C_{K}=(1/2)(3/17)\left(K\left(\bm{0},\bm{0}\right)-K\left(\bm{1},\bm{0}\right)\right)$ for the constant in (12), using (31) along with the fact that $\mathcal{M}^{*}\subset\mathcal{M}_{1}^{+}\left(\mathbb{R}^{d}\right)$, we have that
@@
Consequently the theorem is proved if we can find absolute constants $\alpha_{0},\beta_{0}\in(0,1)$ and pre- and post-change distributions $\mathbb{P},\mathbb{Q}\in\mathcal{M^{*}}$ such that for all $\alpha\leq\alpha_{0}$ it holds that
@@
To show (33), one can use a change of measure argument originally due to Lai [30]. In fact one can directly use the version of Lai’s argument adapted to finite sample analysis by Yu et al. [64, Proposition 4.1]. For clarity of exposition, we repeat the argument below. The following holds for arbitrary $\mathbb{P},\mathbb{Q}\in\mathcal{M^{*}}$. For each $n\in\mathbb{N}$ let $\mathcal{F}_{n}$ be the $\sigma$-field generated by $\{X_{i}\}_{i=1}^{n}$ and let $\mathbb{P}^{\otimes n}$ be the restriction of the joint law to $\mathcal{F}_{n}$. We can write
@@ A.4.5 Proof of Lemma 1
It suffices to show that for any $\epsilon>0$ it holds that
@@
$\mathbb{P}\otimes\Lambda\Bigg(\Bigg\|\frac{1}{n}\sum_{i=1}^{n}\hat{z}_{K}(X_{i})-\mathbb{E}_{\mid\omega}\hat{z}_{K}(X)\Bigg\|>\epsilon\Bigg)\leq 2\exp\Bigg(-\frac{1}{2}\frac{n\varepsilon^{2}}{\tilde{\sigma}^{2}+2\varepsilon}\Bigg),$ (36)
which implies the stated result by using $\pm\mathbb{E}_{\mid\omega}\hat{z}_{K}(X)$, the triangle inequality, and a union bound. Here and in the following $\mathbb{E}_{\mid\omega}$ (resp. $\mathbb{P}_{\mid\omega}$) refers to the expectation (resp. probability) with $\omega$ fixed. To prove (36), we fix $\omega=(\omega_{1},\ldots,\omega_{r})$ and let $\eta_{i}=\hat{z}_{K}(X_{i})-\mathbb{E}_{\mid\omega}\hat{z}_{K}(X)$. Then, all $\eta_{i}$-s are zero mean w.r.t. $\mathbb{P}$ and i.i.d. Further, we have for any $p\geq 2$ that
$\mathbb{E}_{\mid\omega}\|\eta_{1}\|^{p}=\mathbb{E}_{\mid\omega}\|\eta_{1}\|^{p-2}\|\eta_{1}\|^{2}\leq\sup\|\eta_{1}\|^{p-2}\mathbb{E}_{\mid\omega}\|\eta_{1}\|^{2}\leq 2^{p-2}\mathbb{E}_{\mid\omega}\|\eta_{1}\|^{2},$
where the last inequality follows by using the triangle inequality and $\langle\hat{z}_{K}(X_{1}),\hat{z}_{K}(X_{1})\rangle=1$ to obtain
$\sup\|\eta_{1}\|^{p-2}\leq 2^{p-2}\sup\|\hat{z}_{K}(X_{1})\|^{p-2}=2^{p-2}.$
Hence, we have the bound
$\mathbb{E}_{\mid\omega}\|\eta_{1}\|^{p}\leq 2^{p-2}\mathbb{E}_{\mid\omega}\|\eta_{1}\|^{2}\leq\frac{1}{2}p!2^{p-2}\mathbb{E}_{\mid\omega}\|\eta_{1}\|^{2}$
and the $\eta_{i}$-s satisfy the Bernstein condition with $B^{2}=\mathbb{E}_{\mid\omega}\|\eta_{1}\|^{2}$ and $H=2$. The application of Theorem 8 yields that for any $\varepsilon>0$
$\mathbb{P}_{\mid\omega}\Bigg(\Bigg\|\frac{1}{n}\sum_{i=1}^{n}\eta_{i}\Bigg\|>\epsilon\Bigg)\leq 2\exp\Bigg(-\frac{1}{2}\frac{n^{2}\varepsilon^{2}}{B^{2}+\varepsilon nH}\Bigg)\leq 2\exp\Bigg(-\frac{1}{2}\frac{n\varepsilon^{2}}{B^{2}+\varepsilon H}\Bigg).$
To lift the conditioning, we observe that
$\mathbb{P}\otimes\Lambda\Bigg(\Bigg\|\frac{1}{n}\sum_{i=1}^{n}\eta_{i}\Bigg\|>\epsilon\Bigg)$ $=\mathbb{E}\mathbb{P}_{\mid\omega}\Bigg(\Bigg\|\frac{1}{n}\sum_{i=1}^{n}\eta_{i}\Bigg\|>\epsilon\Bigg)\leq 2\mathbb{E}\exp\Bigg(-\frac{1}{2}\frac{n\varepsilon^{2}}{B^{2}+\varepsilon H}\Bigg)$
$\leq 2\exp\Bigg(-\frac{1}{2}\frac{n\varepsilon^{2}}{\mathbb{E}B^{2}+\varepsilon H}\Bigg),$
where the last inequality holds by the concavity of $\exp(-1/x)$ for $x>0$ and Jensen’s inequality. Finally, we conclude the proof by using that
$\mathbb{E}B^{2}$ $=\mathbb{E}\Big(\langle\hat{z}_{K}(X),\hat{z}_{K}(X)\rangle+\langle\mathbb{E}_{\mathbb{P}}\hat{z}_{K}(X),\mathbb{E}_{\mathbb{P}}\hat{z}_{K}(Y)\rangle-2\langle\hat{z}_{K}(X),\mathbb{E}_{\mathbb{P}}\hat{z}_{K}(Y)\Big)$
$=\mathbb{E}K(X,X)-\mathbb{E}K(X,Y)=\tilde{\sigma}^{2},$
where the boundedness of all terms allows exchanging the expectations by the Fubini-Tonelli theorem. ∎
A.4.6 Proof of Theorem 5
Proof.
We first show that with probability at least $1-\alpha$ no local tests conducted on null intervals will cause the algorithm to wrongly declare a change. For each $n\leq T$ define $k_{n}=\max\{k\leq M+1\mid\eta_{k}\leq n\}$. With $\pi_{n}$ defined as in (17), introduce the event
$A_{1}=\left\{\cap_{n=2}^{T}\cap_{j=1}^{\left\lfloor\log_{2}(n-\eta_{k_{n}}+1)\right\rfloor-1}\hat{M}_{n,j}\leq\pi_{n}\right\}.$ (37)
We therefore have that
$\mathbb{P}(A_{1}^{c})$ $\leq\sum_{l=1}^{T}2^{l}\max_{2^{l}\leq n\leq 2^{l+1}}\sum_{j=0}^{\left\lfloor\log_{2}(n-\eta_{k_{n}}+1)\right\rfloor-1}\mathbb{P}\left(\hat{M}_{n,j}>\pi_{n}\right)$ (38a)
$\leq\sum_{l=1}^{T}l2^{l}e^{-(\pi_{2^{l}})^{2}/2}$ (38b)
$=\frac{\alpha\log_{2}(T)}{1+\log_{2}(T)}<\alpha,$
where (38a) follows by arguments similar to (18a) and (38b) holds due to Lemma 5. Therefore, if $M=0$ we are done. We now tackle the case $M>0$. We will show that when $M>0$ with probability at least $1-2\alpha$ all true change points are detected and localized at the stated rate. We first make the constants in Theorem 5 explicit:
$C_{1}=2\times C_{3}$
$C_{2}=4$
$C_{3}=\left(2\times\left(\sqrt{2}+\sqrt{4}+\sqrt{6}\right)\right)^{2}.$
Introduce also the event